OpenStudy (anonymous):

Factor completely: 2x^2 + 2x - 12?

OpenStudy (anonymous):

A. 2(x-3)(x+4) B. 2(x-2)(x+3) C.(2x-4)(x+3) D.(2x -3)(x+4)

OpenStudy (praxer):

2x^2+2x-12=0 =>x^2+x-6=0 =>x^2+3x-2x-6=0 =>(x-2)(x+3) now multiply it by 2 we get 2(x-2)(x+3). So,option B. is correct.... :)

OpenStudy (anonymous):

Awesome i thought it was B too :) thank you!!

OpenStudy (praxer):

$$\huge \frak Welcome \ddot\smile$$